### v1

@@ Appendix A Calibration of MUSE catalogues
We downloaded the PSF-extracted HST catalogues with transformed magnitudes in the Johnson’s system171717Catalogues can be found at http://astronomy.nmsu.edu/holtz/archival/ic1613/html/ic1613.html. The covered photometric bands were V and I, with the addition of the B-band for the central HST pointing. The Subaru catalogue covered instead the Johnson’s B and V bands. These are deep photometric catalogues, reaching down to the horizontal branch (at ), which means that they cover a much wider photometric range than that of our MUSE data. Before proceeding, we performed a pre-cleaning, selecting all sources marked as stellar in the HST catalogues, while retaining those targets in the Subaru catalogue having DAOPHOT parameters sharpness (SHARP) between and , and a goodness-of-fit (CHI) .
@@ B.1 Verify SPEXXY general performance
We verified the performance of spexxy in recovering l.o.s. velocities and associated errors. We recall that spexxy performs a full spectral fit to the observed spectra using interpolated templates generated from the PHOENIX library of high-resolution synthetic spectra (Husser et al. 2013). It requires a configuration file specifying initial values for velocity, effective temperature, surface gravity, metallicity and [/M] from which to start the fit. We obtained the initial values from a set of isochrones (Girardi et al. 2000; Bressan et al. 2012) with dex (around the mean metallicity of IC 1613) by performing a linear regression between the observed colours and magnitudes and the theoretical and . For the spectral fitting, we used the available PHOENIX templates, convolved with the MUSE line spread function, with solar scale chemical composition, thus fixing dex. We let the code find the best fit of all other parameters (i.e., , and , together with the l.o.s. velocity). We also masked certain spectral regions to avoid possible emission lines due to the interstellar medium and residuals due to sky line subtraction (such as the persistent residual due to the [OI] line at 5577 Å). To test for possible systematic velocity shifts we ran the code on several noiseless templates from the PHOENIX library of giant stars at different (from 2500 K to 15000 K) at fixed dex. Templates were shifted at several l.o.s. velocities, from km s to km s at step of 50 km s. As initial guess for the velocity fitting, we set km s (i.e. approximately IC 1613 systemic velocity). We found that in general spexxy correctly recovers the assigned l.o.s. velocities (mean and scatter of the difference between input and output velocities being km s), although we noticed some failure around the extremes (i.e. for km s and km s) with velocities variations of up to 50%. This occurred in only 3% of the cases that are those differing the most from the initial guess for the velocity. We recall that based on the Besançon model we expect almost no contaminants with true velocities km s, while the variation at km s is small enough to keep such possible contaminants away from the bulk velocity of IC 1613.
@@
Since we have two main stellar populations in our observed sample, RGS and MS stars, we generated mock spectra representative of these two sub-samples. For the RGS, we selected stars with K, with a difference between observed values found by spexxy and ULySS of no more than 500 K. From their best-fitting templates and scaled noise spectra, we generated 250 mock spectra in bins of S/N around . The same was done for the MS stars, but for stars with K, generating mock spectra at S/N around . Results are reported in Table 6.
We found that for the RGS stars their velocities and associated errors are well recovered, with deviations between the mean of the errors and the measurement scatter within km s. The effective temperature was also well recovered, although the errors were as much as K lower than the measurement scatter. On the other hand, and were relatively more uncertain, with errors underestimated between 0.1 and 0.2 dex (higher deviations at lower S/N), leading to an offset on the recovered values of up to 0.1 dex.
For the MS stars we recovered higher deviations from the input values than the RGS case. In particular, velocity errors at S/N were underestimated by more than km s, while at the highest S/N the deviation was less than 1 km s. In the main text we reported that we made a quality cut in our sample by excluding stars with observed velocity errors km s. This cut mainly excluded MS stars with the lowest S/N, while ensuring that we kept those with well-estimated velocity errors within 10%. For the spectral parameters, the associated errors were generally more underestimated () than in the RGS case.
@@ B.2 Comparison between SPEXXY and ULySS
In our analysis, we used spexxy and ULySS in a complementary way: the former for l.o.s. velocity measurements, while the latter for spectral classification. However, since both codes provide similar outputs, we compared their performances in recovering the l.o.s. velocity and from the observed spectra. We considered all targets in the sample used for the kinematic analysis (see Sect. 4). The l.o.s. velocity errors for the ULySS measurements were obtained from Monte Carlo simulations using several templates of giant stars at different intervals of temperature and S/N. We made this choice because the formal errors returned by ULySS were generally overestimated, without following any clear trend with the S/N or spectral type. For the Monte Carlo simulations we chose seven templates from the MIUSCAT library (see Sect.3.1), with ranging from 3500 K to 20000 K. The simulations were performed by adding Poisson noise to each template to obtain simulated spectra at increasing S/N intervals, with values ranging from 4 to 30. We repeated each process 100 times. For each template, the recovered velocity scatter as a function of S/N was fitted by an exponential law. The ULySS measurements were then divided into effective temperature bins according to the templates considered, and the velocity errors assigned using the corresponding exponential fit. First we compared the velocity differences for pair measurements obtained with ULySS and spexxy. We found that % of the measured velocities are in agreement within the combined uncertainties, obtained as . The comparison is shown in the left panel of Fig. 10 where for clarity we reported only those measurements with a combined error km s. Nevertheless, the difference between the ULySS and spexxy measurements showed an average positive shift of km s. What is more, a higher shift of km s is observed for those objects with a S/N. This was somehow expected for the UlySS measurements, since the Monte Carlo tests showed average velocity shifts of up to 20 km s for templates with effective temperatures of K and greater than K having simulated . On the other hand, we have already showed in the previous section the capacity of spexxy to correctly recover l.o.s. velocities, even at the lowest S/N for red giant stars. This is a very important point because most of the stars in our dataset are RGS in a low S/N regime (), where spexxy showed its strength in the recovery of l.o.s. velocities. The comparison of effective temperatures, as shown in the right-hand panel of Fig. 10, gave good results for K, although the spexxy values tend to be systematically K higher than the ULySS values. At higher , spexxy values started to deviate significantly. This is mainly due to the fact that spexxy measurements are based on the PHOENIX library, which is not optimized for very hot stars. In the ULySS case, instead, for each spectrum the stored temperature value is that of the best fitting template, which was chosen by eye among the most likely templates. We recall that this procedure had an accuracy of up to K, for a quality PCL-flag between 3 and 4.

### v2

@@ Appendix A Calibration of MUSE catalogues
We downloaded the PSF-extracted HST catalogues with transformed magnitudes in the Johnson’s system191919Catalogues can be found at http://astronomy.nmsu.edu/holtz/archival/ic1613/html/ic1613.html. The covered photometric bands were V and I, with the addition of the B-band for the central HST pointing. The Subaru catalogue covered instead the Johnson’s B and V bands. These are deep photometric catalogues, reaching down to the horizontal branch (at ), which means that they cover a much wider photometric range than that of our MUSE data. Before proceeding, we performed a pre-cleaning, selecting all sources marked as stellar in the HST catalogues, while retaining those targets in the Subaru catalogue having DAOPHOT parameters sharpness (SHARP) between and , and a goodness-of-fit (CHI) .
@@ B.1 Verify SPEXXY general performance
We verified the performance of spexxy in recovering l.o.s. velocities and associated errors. We recall that spexxy performs a full spectral fit to the observed spectra using interpolated templates generated from the PHOENIX library of high-resolution synthetic spectra (Husser et al. 2013). It requires a configuration file specifying initial values for velocity, effective temperature, surface gravity, metallicity and [/M] from which to start the fit. We obtained the initial stellar parameters from a set of isochrones (Girardi et al. 2000; Bressan et al. 2012) with dex (around the mean metallicity of IC 1613) by performing a linear regression between the observed colours and magnitudes and the theoretical and . We also set an initial velocity guess of km s-1 (i.e. approximately IC 1613 systemic velocity). For the spectral fitting, we used the available PHOENIX templates, convolved with the MUSE line spread function, with solar scale chemical composition, thus fixing dex. We let the code find the best fit of all other parameters (i.e., , and , together with the l.o.s. velocity). We also masked certain spectral regions to avoid possible emission lines due to the interstellar medium and residuals due to sky line subtraction (such as the persistent residual due to the [OI] line at 5577 Å).
To test for possible systematic velocity shifts we ran the code on several noiseless templates from the PHOENIX library of giant stars at different (from 2500 K to 15000 K) at fixed dex. Templates were shifted at several l.o.s. velocities, from km s-1 to km s-1 at step of 50 km s-1. As initial guess for the velocity fitting, we set again km s-1. We found that in general spexxy correctly recovers the assigned l.o.s. velocities (mean and scatter of the difference between input and output velocities being km s-1), although we noticed some failure around the extremes (i.e. for km s-1 and km s-1) with velocities variations of up to 50%. This occurred in only 3% of the cases that are those differing the most from the initial guess for the velocity. We recall that based on the Besançon model we expect almost no contaminants with true velocities km s-1, while the variation at km s-1 is small enough to keep such possible contaminants away from the bulk velocity of IC 1613.
@@
Since we have two main stellar populations in our observed sample, RGS and MS stars, we generated mock spectra representative of these two sub-samples. For the RGS, we selected stars with K, with a difference between observed values found by spexxy and ULySS of no more than 500 K. From their best-fitting templates and scaled noise spectra, we generated 250 mock spectra in bins of S/NCaT around . The same was done for the MS stars, but for stars with K, generating mock spectra at S/N550 around . Results are reported in Table 6.
We found that for the RGS stars their velocities and associated errors are well recovered, with deviations between the mean of the errors and the measurement scatter within km s-1. The effective temperature was also well recovered, although the errors were as much as K lower than the measurement scatter. On the other hand, and were relatively more uncertain, with errors underestimated between 0.1 and 0.2 dex (higher deviations at lower S/N), leading to an offset on the recovered values of up to 0.1 dex.
For the MS stars we recovered higher deviations from the input values than the RGS case. In particular, velocity errors at S/N were underestimated by more than km s-1, while at the highest S/N the deviation was less than 1 km s-1. In the main text we reported that we made a quality cut in our sample by excluding stars with observed velocity errors km s-1. This cut mainly excluded MS stars with the lowest S/N, while ensuring that we kept those with well-estimated velocity errors within 10%. For the spectral parameters, the associated errors were generally more underestimated () than in the RGS case.
@@ B.2 Comparison between SPEXXY and ULySS
In our analysis, we used spexxy and ULySS in a complementary way: the former for l.o.s. velocity measurements, while the latter for spectral classification. However, since both codes provide similar outputs, we compared their performances in recovering the l.o.s. velocity and from the observed spectra.
@@
We considered all targets in the sample used for the kinematic analysis (see Sect. 4). The l.o.s. velocity errors for the ULySS measurements were obtained from Monte Carlo simulations using several templates of giant stars at different intervals of temperature and S/N. We made this choice because the formal errors returned by ULySS were generally overestimated, without following any clear trend with the S/N or spectral type. For the Monte Carlo simulations we chose seven templates from the MIUSCAT library (see Sect.3.1.1), with ranging from 3500 K to 20000 K. The simulations were performed by adding Poisson noise to each template to obtain simulated spectra at increasing S/N intervals, with values ranging from 4 to 30. We repeated each process 100 times. For each template, the recovered velocity scatter as a function of S/N was fitted by an exponential law. The ULySS measurements were then divided into effective temperature bins according to the templates considered, and the velocity errors assigned using the corresponding exponential fit.
First we compared the velocity differences for pair measurements obtained with ULySS and spexxy. We found that % of the measured velocities are in agreement within the combined uncertainties, obtained as . The comparison is shown in the left panel of Fig. 11 where for clarity we reported only those measurements with a combined error km s-1. Nevertheless, the difference between the ULySS and spexxy measurements showed an average positive shift of km s-1. What is more, a higher shift of km s-1 is observed for those objects with a S/N. This was somehow expected for the UlySS measurements, since the Monte Carlo tests showed average velocity shifts of up to 20 km s-1 for templates with effective temperatures of K and greater than K having simulated . On the other hand, we have already showed in the previous section the capacity of spexxy to correctly recover l.o.s. velocities, even at the lowest S/N for red giant stars. This is a very important point because most of the stars in our dataset are RGS in a low S/N regime (), where spexxy showed its strength in the recovery of l.o.s. velocities.
The comparison of effective temperatures, as shown in the right-hand panel of Fig. 11, gave good results for K, although the spexxy values tend to be systematically K higher than the ULySS values. At higher , spexxy values started to deviate significantly. This is mainly due to the fact that spexxy measurements are based on the PHOENIX library, which is not optimized for very hot stars. In the ULySS case, instead, for each spectrum the stored temperature value is that of the best fitting template, which was chosen by eye among the most likely templates. We recall that this procedure had an accuracy of up to K, for a quality PCL-flag between 3 and 4.
B.3 Biases in the recovery of kinematic properties
We checked for the presence of possible biases that could alter the recovery of the kinematic properties of the stellar samples, in particular when inspecting the properties of the RGS and MS stars. On one hand, the region approximately separating the MS from the RGS, (see Fig. 4), contains the instability strip of IC 1613, where it is expected to find mainly classical Cepheids with short-term variability (i.e., with log and Gyr; Udalski et al. 2001; Bernard et al. 2010, see e.g.). Since we observe them in random phases, we expect any differences to be averaged out. It is also true that, especially for Cepheids, hydrogen or metallic lines can lead to different velocities (see, e.g., Vinko et al. 1998). Nonetheless, we find no significant differences when calculating the average velocities and velocity dispersion including or excluding stars in that color range. Similar considerations can be made for the Be-star candidates in our sample, which should also be short-period variables (e.g., Porter & Rivinius 2003; Rivinius et al. 2013), although in this case it is mostly the intensity of the emission lines that is affected. We repeated the same test finding values compatible within the errors. As for the RGS sample, this contains a small number of C-star candidates, which are long period variables (e.g., Menzies et al. 2015). However, we have verified that the C-stars have similar properties to the rest of the sample within the errors.
We did not check for the presence of binary stars, which could particularly affect the velocity dispersion values. The low S/N of the individual exposures for each field in our sample and the short period over which they were observed (between a few days and a month), preclude a quantitative analysis. We can, however, make qualitative estimates. For a sample of mostly RGB stars with high measured velocity dispersion ( km s-1), as it is our case, the impact of unresolved binaries is expected to be minimal (i.e. of inflation; Minor et al. 2010; Minor 2013; Spencer et al. 2017, 2018; Arroyo-Polonio et al. 2023, see). However, the case is different when focusing on the MS stars sample. The multiplicity fraction (i.e. the fraction of stars having at least one companion) is generally expected to be for A-types, which increase to for OB stars (e.g. Moe & Di Stefano 2017; Offner et al. 2023). In particular, these stars are often found in triples or higher order multiples. For Be stars such figures are less constrained, but mass accretion through binary interactions has been proposed as a possible channel for their formation (Pols et al. 1991; de Mink et al. 2013; Shao & Li 2014; Bodensteiner et al. 2020).
Considering that the multiple OBs cover a range of orbital periods, preferably a few days or a few decades (Moe & Di Stefano 2017), the stacking process to generate the spectra should already average the short-period ones. For the long-period ones (i.e. log) we expect velocity variations of km s-1, which could translate into a velocity dispersion inflation of (following Minor et al. 2010, assuming that they behave as short-period low-mass binaries and have intrinsic km s-1). This would especially affect the central pointing, where the fraction of OB stars is high (). It could thus partly explain why in this field the value of the velocity dispersion of the MS stars is comparable to that of the RGS, while in the other fields it tends to be closer to that of the HI. Another possible source of disturbance for the MS stars in F1 is that they may still be coupled to the motion of the ionised bubble, which is probably expanding (with a typical velocity of km s-1, Pokhrel et al. 2020). We also do not exclude the presence of possible runaway OB stars that could further impact the kinematics of the sample (e.g. Stoop et al. 2024).